Morgana Company identifies three activities in its manufacturing process: machine setups, machining, and inspections. Estimated annual overhead cost for each activity is $201,600, $413,100, and $84,600, respectively. The cost driver for each activity and the estimated annual usage are number of setups 2,800, machine hours 24,300, and number of inspections 1,800.

Compute the overhead rate for each activity.

Machine setups $ per setup
Machining $ per machine hour
Inspections $ per inspection
Answer #1

The overhead rate is a cost allocated to the production of a product or service. To allocate overhead costs, an overhead rate is applied to the direct costs tied to production by spreading or allocating the overhead costs based on specific measures.

 Overhead Rate = Costs/Allocation Measure. 

a.Overhead rate for Machine Setups

Costs = $201,600

Number of Setups=2800

Overhead rate= 201600/2800= $72 per setup

b.Overhead rate for Machining

Costs = $413,100

Total Machine Hours=24,300

Overhead rate=413100/24300= $17 per hour

c.Overhead rate for Inspections

Costs= $84600

Number of inspections= 1800

Overhead rate=84600/1800=$47 per inspection


answered by: gavin
Answer #2

Overhead Rate

The overhead rate is a cost allocated to the production of a product or service. To allocate overhead costs, an overhead rate is applied to the direct costs tied to production by spreading or allocating the overhead costs based on specific measures.

Overhead Rate = Costs/Allocation Measure.

a.Overhead rate for Machine Setups

Costs = $201,600

Number of Setups=2800

Overhead rate= 201600/2800= $72 per setup

b.Overhead rate for Machining

Costs = $413,100

Total Machine Hours=24,300

Overhead rate=413100/24300= $17 per hour

c.Overhead rate for Inspections

Costs= $84600

Number of inspections= 1800

Overhead rate=84600/1800=$47 per inspection